# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.1.2 Properties of the value function
$\begin{cases}\partial_{t}w=\frac{1}{2}\varepsilon^{2}D^{2}_{xx}w+\frac{1}{2}\chi^{2}D^{2}_{yy}w-\frac{1}{2}\chi D_{y}w+\rho w+H_{0}(D_{x}w)+H_{1}(T-t,x,e^{y},\mu(T-t),e^{-y}D_{y}w)&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}\\ w(0,x,y)=\phi(x,e^{y})\quad\mbox{ in }\mathbb{R}\times\mathbb{R}\\ w(t,x,-\infty)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (27)
@@
$\mathcal{L}\psi:=\frac{1}{2}\varepsilon^{2}D^{2}_{xx}\psi+\frac{1}{2}\chi^{2}D^{2}_{yy}\psi-\frac{1}{2}\chi D_{y}w+\rho\psi$
@@
from which the claim follow setting
@@
The function $D_{p}H_{1}$ is locally Lispchitz continuous for $M(\mu)$ away from 00, in the following sense: for every $N\geq 1$ there exists a positive constant $C_{N}$ such that for every $\mu_{1},\mu_{2}$ satisfying $\frac{1}{N}<M(\mu_{1})\wedge M(\mu_{2})\leq M(\mu_{1})\vee M(\mu_{2})<N$, for every $h_{1}\leq h_{2}<N$, for every $x_{1},x_{2}$ and for every $p_{1},p_{2}$ we have
@@
otherwise, so that $p>p_{0}(x,h,\mu)$ if and only if $F(x,\mu)>F_{0}(x,h,p)$, and we have $F_{0}\left(x,h,p_{0}(x,h,\mu)\right)=F(x,\mu)$ whenever $p$ and $h$ are nonzero. Consider $\mu_{1},\mu_{2}$ such that $0<\underline{M}:=M(\mu_{1})\wedge M(\mu_{2})$, and set also $\overline{M}:=M(\mu_{1})\vee M(\mu_{2})$, $\alpha=\frac{1-\gamma(1-\sigma)}{1-\eta}$ and $\beta=\frac{1-\sigma}{1-\eta}$. If $F_{0}(x,h,p)>F(x,\mu_{1})\vee F(x,\mu_{2})$ then
@@
$\leq f(h)\frac{\Theta}{\theta^{2}}\left(\frac{M(\mu_{1})}{L_{\eta_{2}}}+\Theta\right)\mathcal{W}_{1}(\mu_{1},\mu_{2})+p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\left|F(x,\mu_{1})-F(x,\mu_{2})\right|$
$\leq f(h)\frac{\Theta}{\theta^{2}}\left(\frac{M(\mu_{1})}{L_{\eta_{2}}}+\Theta\right)\mathcal{W}_{1}(\mu_{1},\mu_{2})\left(1+F(x,\mu_{1})^{\alpha}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\right)$
$\leq L_{f}h\frac{\Theta}{\theta^{2}}\left(\frac{\overline{M}}{L_{\eta_{2}}}+\Theta\right)\mathcal{W}_{1}(\mu_{1},\mu_{2})\left(1+\left|1-\alpha\right|\left(\frac{\Theta^{2}}{\theta^{2}}\frac{\overline{M}}{\underline{M}}\right)^{\alpha}\right)$
@@
$\leq\frac{\Theta}{\theta}M(\mu_{1})f(h)F(x,\mu_{1})^{\alpha}\left(1-\gamma(1-\sigma)\right)\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\frac{\Theta}{\theta^{2}}\left(\frac{M(\mu_{1})}{L_{\eta_{2}}}+\Theta\right)\mathcal{W}_{1}(\mu_{1},\mu_{2})$
$\leq\left(\frac{\Theta}{\theta}\overline{M}\right)^{1+\alpha}L_{f}h(1-\gamma(1-\sigma))\left(\frac{\Theta}{\theta}\frac{1}{\underline{M}}\right)^{2-\gamma(1-\sigma)}\frac{\Theta}{\theta^{2}}\left(\frac{M(\mu_{1})}{L_{\eta_{2}}}+\Theta\right)\mathcal{W}_{1}(\mu_{1},\mu_{2}).$
@@
The result follows substituting, in all bounds found above, $\underline{M}$ with $\frac{1}{N}$, $\overline{M}$ with $N$, $h$ with $N$ and $M(\mu)$ with $\frac{1}{N}$ if it is raised to a negative exponent and with $N$ otherwise. ∎
@@
$\begin{cases}\partial_{t}\mu(t)=\frac{1}{2}\varepsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0}\end{cases}$ (34)
@@
$\mathcal{Q}_{K_{1},K_{2}}=\left\{\mu(\cdot)\colon[0,T]\rightarrow\mathcal{P}_{1}\text{ s.t. }\mu(0)=\mu_{0},\sup_{t\in[0,T]}P_{2}(\mu(t))\leq 2K_{1},\sup_{\begin{subarray}{c}s,t\in[0,T]\\ s\neq t\end{subarray}}\frac{\mathcal{W}_{2}(\mu(s),\mu(t))}{|s-t|^{\frac{1}{2}}}\leq K_{2}\right\}.$ (39)
@@ 4 Solution to the Mean Field Game
$C(\mu,\nu,\mu_{0},T)=24T^{2}\left(\frac{\Theta}{\theta^{2}}\frac{1}{L_{\eta_{2}}}\max\left\{P(\mu),P(\nu)\right\}+\frac{\Theta^{2}}{\theta^{2}}\right)^{2}\mathbb{E}\left[h_{0}^{2}\right]e^{TC_{2,2}(T,\nu(\cdot))+6TL_{f}^{2}\frac{\Theta^{2}}{\theta^{2}}M(\mu)^{2}+3T\zeta^{2}+3\chi^{2}}.$ (48)
@@
[…]

### v2

@@ 3.1.2 Properties of the value function
$\begin{cases}\partial_{t}w=\frac{1}{2}\epsilon^{2}D^{2}_{xx}w+\frac{1}{2}\chi^{2}D^{2}_{yy}w-\frac{1}{2}\chi D_{y}w+\rho w+H_{0}(D_{x}w)+H_{1}(T-t,x,e^{y},\mu(T-t),e^{-y}D_{y}w)&\mbox{ in }(0,T)\times\mathbb{R}\times\mathbb{R}\\ w(0,x,y)=\phi(x,e^{y})\quad\mbox{ in }\mathbb{R}\times\mathbb{R}\\ w(t,x,-\infty)=0\quad\mbox{ in }(0,T)\times\mathbb{R}\end{cases}$ (27)
@@
$\mathcal{L}\psi:=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\psi+\frac{1}{2}\chi^{2}D^{2}_{yy}\psi-\frac{1}{2}\chi D_{y}w+\rho\psi$
@@
from which the claim follows setting
@@
The function $D_{p}H_{1}$ is locally Lispchitz continuous for $M(\mu)$ away from 00, in the following sense: for every $N\geq 1$ there exists a positive constant $C_{N}$ such that for every $\mu_{1},\mu_{2}$ satisfying $\frac{1}{N}<M(\mu_{1})\wedge M(\mu_{2})\leq\left(M_{2}(\mu_{1})\vee M_{2}(\mu_{2})\right)^{\frac{1}{2}}<N$, for every $h_{1}\leq h_{2}<N$, for every $x_{1},x_{2}$ and for every $p_{1},p_{2}$ we have
@@
otherwise, so that $p>p_{0}(x,h,\mu)$ if and only if $F(x,\mu)>F_{0}(x,h,p)$, and we have $F_{0}\left(x,h,p_{0}(x,h,\mu)\right)=F(x,\mu)$ whenever $p$ and $h$ are nonzero. Consider $\mu_{1},\mu_{2}$ such that $0<\underline{M}:=M_{2}(\mu_{1})\wedge M_{2}(\mu_{2})$, and set also $\overline{M}_{2}:=M_{2}(\mu_{1})\vee M_{2}(\mu_{2})$, $\alpha=\frac{1-\gamma(1-\sigma)}{1-\eta}$ and $\beta=\frac{1-\sigma}{1-\eta}$. If $F_{0}(x,h,p)>F(x,\mu_{1})\vee F(x,\mu_{2})$ then
@@
$\leq f(h)\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})$
$\phantom{aaaaaaaaaa}+p_{0}(x,h,\mu_{1})^{-\frac{1}{1-\eta}}(1-\gamma)^{\frac{1}{1-\eta}}A(x)^{\beta}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\left|F(x,\mu_{1})-F(x,\mu_{2})\right|$
$\leq f(h)\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})\left(1+F(x,\mu_{1})^{\alpha}\left|1-\alpha\right|\left(\frac{\theta}{\Theta}\underline{M}\right)^{-\alpha}\right)$
$\leq L_{f}h\frac{\Theta}{\theta^{2}}\left(\overline{M}_{2}^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})\left(1+\left|1-\alpha\right|\left(\frac{\Theta^{2}}{\theta^{2}}\frac{\overline{M}_{2}^{\frac{1}{2}}}{\underline{M}}\right)^{\alpha}\right)$
@@
$\leq\frac{\Theta}{\theta}M(\mu_{1})f(h)F(x,\mu_{1})^{\alpha}\left(1-\gamma(1-\sigma)\right)\left(\frac{\theta}{\Theta}\underline{M}\right)^{\gamma(1-\sigma)-2}\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2})$
$\leq\left(\frac{\Theta}{\theta}\overline{M}_{2}^{\frac{1}{2}}\right)^{1+\alpha}L_{f}h(1-\gamma(1-\sigma))\left(\frac{\Theta}{\theta}\frac{1}{\underline{M}}\right)^{2-\gamma(1-\sigma)}\frac{\Theta}{\theta^{2}}\left(M_{2}(\mu_{1})^{\frac{1}{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}\right)+\Theta\right)\mathcal{W}_{2}(\mu_{1},\mu_{2}).$
@@
The result follows substituting, in all bounds found above, $\underline{M}$ with $\frac{1}{N}$, $\overline{M}_{2}^{\frac{1}{2}}$ with $N$, $h$ with $N$ and $M(\mu)$ with $\frac{1}{N}$ if it is raised to a negative exponent and with $N$ otherwise. ∎
@@
$\begin{cases}\partial_{t}\mu(t)=\frac{1}{2}\epsilon^{2}D^{2}_{xx}\mu(t)+\frac{1}{2}\chi^{2}D^{2}_{hh}\left(h^{2}\mu(t)\right)-D_{x}\left(D_{p}H_{0}(D_{x}V(t,x,h))\mu(t)\right)-D_{h}\left(D_{p}H_{1}(x,h,\mu(t),D_{h}V(t,x,h))\mu(t)\right),\\ \mu(0)=\mu_{0}\end{cases}$ (34)
@@
$\mathcal{Q}_{K_{1},K_{2}}=\left\{\mu(\cdot)\colon[0,T]\rightarrow\mathcal{P}_{2}\text{ s.t. }\mu(0)=\mu_{0},\sup_{t\in[0,T]}P_{2}(\mu(t))\leq 2K_{1},\sup_{\begin{subarray}{c}s,t\in[0,T]\\ s\neq t\end{subarray}}\frac{\mathcal{W}_{2}(\mu(s),\mu(t))}{|s-t|^{\frac{1}{2}}}\leq K_{2}\right\}.$ (39)
@@ 4 Solution to the Mean Field Game
$C(\mu,\nu,\mu_{0},T)=24T^{2}\left(\frac{\Theta}{\theta^{2}}\left(L_{\eta_{1}}+L_{\eta_{2}}^{-1}\right)\max\left\{P_{2}(\mu)^{\frac{1}{2}},P_{2}(\nu)^{\frac{1}{2}}\right\}+\frac{\Theta^{2}}{\theta^{2}}\right)^{2}\mathbb{E}\left[h_{0}^{2}\right]e^{TC_{2,2}(T,\nu(\cdot))+6TL_{f}^{2}\frac{\Theta^{2}}{\theta^{2}}M(\mu)^{2}+3T\zeta^{2}+3\chi^{2}}.$ (48)
@@
By the regularity of the value function for each $\mu_{n}(\cdot)$, we know that $V_{n}$ is a solution of the Hamilton-Jacobi-Bellman equation in (10) where $H_{1}$ is evaluated in $\mu_{n}(\cdot)$. Then, by a similar procedure as the one in the proof of Theorem 4.4, by Lemma 4.2 and by the Ascoli-Arzelà theorem coupled with a diagonal argument, one could directly infer that $V_{n}$ converges locally uniformly to some $v$ and by stability of viscosity solutions applied to the HJB equation, one could then deduce that the subsequence $V_{n}$ converges to a solution $v$ of the same equation with $\tilde{H}_{1}$ evaluated at $\mu(\cdot)$. In any case, since no uniqueness of the solution of the HJB equation is guaranteed, it is not possible to conclude that the function $v$ obtained in this way is the value function corresponding to $\mu$. Therefore, due to how the operator $\Psi_{\mu_{0}}$ is defined, in the previous proof we need to prove directly the convergence of $V_{n}$ to $V$.